OpenStudy (anonymous):

The general solution of the differential equation dy − 0.2x dx = 0 is a family of curves. These curves are all lines hyperbolas parabolas ellipses

OpenStudy (anonymous):

i'm thinking hyperbolas but not sure

OpenStudy (anonymous):

\[dy=0.2x dx\] by moving over \[\int\limits dy=\int\limits 0.2x dx\] Is that what you did?

OpenStudy (anonymous):

and then I got y=0.2 which is a graph of a straight line at y=0.2, so I would call it a line.

OpenStudy (anonymous):

um .. no i just shamelessly plugged it into wolfram :P

OpenStudy (anonymous):

Does this involve the slope field?

OpenStudy (anonymous):

it might, some of the other questions iv been asked have involved the slope field

OpenStudy (anonymous):

plug the equation into wolfram and scroll down to the family curve , it seems to show a parabola or hyperbola

OpenStudy (michele_laino):

after a simple integration I got this: \[\Large \int {dy = 0.2\int {xdx} = 0.2 \cdot \frac{{{x^2}}}{2}} + C\]

OpenStudy (anonymous):

Seems like a parabola. Any 2 curves symmetrical is a hyperbola.

OpenStudy (anonymous):

oh yea. I took derivative instead of integrating it..

OpenStudy (anonymous):

okay can you just explain to me what the hell family curves are ?

OpenStudy (michele_laino):

since C is an arbitrary real constant, whose values can vary inside the set of real numbers, in other words, we have: \[\Large C \in \mathbb{R}\]

OpenStudy (michele_laino):